OpenStudy (anonymous):

The specific gravity of a solution of KOH at 15C is 1.0824 and contains 0.813 lb KOH per gal of solution. What are the mass fractions of KOH and H2O in the solution?

OpenStudy (raspberryjam):

Specific gravity is a ratio of densities, in this case it should be \[\frac{density of KOH}{density of water}\] So given that the solution contains .813 lb/gal of KOH... \[\frac{8.13 lb/gallons}{density of water}=1.0824\] Or you can just write it as a function where you have to solve for x: \[\frac{8.13}{x}=1.0824\] So then just solve for the bottom, but remember the units. When solving for x like this you get x in pounds per gallon. Not really sure if they are looking for grams per milliliter, if they are then you simply have to convert the units.